# Mixed Signals: A Diverse Point Cloud Dataset for Heterogeneous LiDAR V2X Collaboration

### v1

@@ Abstract
Vehicle-to-everything (V2X) collaborative perception has emerged as a promising solution to address the limitations of single-vehicle perception systems. However, existing V2X datasets are limited in scope, diversity, and quality. To address these gaps, we present Mixed Signals, a comprehensive V2X dataset featuring 45.1k point clouds and 240.6k bounding boxes collected from three connected autonomous vehicles (CAVs) equipped with two different types of LiDAR sensors, plus a roadside unit with dual LiDARs. Our dataset provides precisely aligned point clouds and bounding box annotations across 10 classes, ensuring reliable data for perception training. We provide detailed statistical analysis on the quality of our dataset and extensively benchmark existing V2X methods on it. Mixed Signals is ready-to-use, making it one of the highest quality, large-scale datasets publicly available for V2X perception research. Details on the website: https://mixedsignalsdataset.cs.cornell.edu/.
Figure 1: Six samples of the scenes in the dataset. The different agent’s LiDAR point clouds are colored according to electric vehicle-001 (EV-1) in purple, electric vehicle-002 (EV-2) in red, urban vehicle (Laser) in yellow, and the RSU DOME and TOP LiDARs are colored in green and blue, respectively. We also draw the annotated bounding boxes within the scene. Best viewed in colour.
@@ 1 Introduction
In recent years, driver assistance [19, 29] and autonomous driving [46, 1] technologies have advanced significantly, equipping vehicles with promising capabilities in perception [20, 43], planning [13, 15], and control [2, 8]. Most of these developments, however, focus on scenarios involving a single autonomous vehicle, where challenges remain in complex or unpredictable situations [41]. For instance, important traffic participants can be occluded from view, or sensors can fail unexpectedly. With more autonomous vehicles being deployed, new possibilities emerge to address these issues: Multiple vehicles can communicate with each other and with nearby infrastructure so that each vehicle can use shared information to reliably detect road users even when its own sensors miss them. This approach is commonly referred as vehicle-to-everything (V2X) collaborative perception.
While single-vehicle perception datasets are abundant and cover diverse driving conditions [11, 18, 33, 4, 5, 28, 9, 38, 16, 27, 30], real-world V2X datasets remain limited in availability, diversity, and quality. Only a handful of publicly available V2X datasets exist [44, 45, 42, 37, 47], with some of them accessible only within specific geographical regions [44, 45]. These data are collected exclusively from three right-hand driving locations, overlooking the unique traffic dynamics in left-hand driving countries which make up about a third of the world [39]. Furthermore, in these datasets, the connected autonomous vehicles (CAVs) share identical LiDAR sensor setups, under the assumption that the communicating vehicles are operated by the same company. However, as collaborative perception becomes more widespread, it would be valuable for vehicles of different models and manufacturers to communicate. These vehicles are likely equipped with different sensor configurations, introducing additional complexities that need to be considered. Finally, as the V2X setting involves multiple agents and sensors, data collection and alignment present additional challenges. Often times, difficulty with pose estimation and faulty localization systems result in poor alignment (Figure 4). Such inaccuracies can lead to suboptimal performance for detector training [40].
To address current limitations, we introduce the Mixed Signals dataset, designed to support diverse real-world V2X research scenarios with clean, high-quality data. Notably, Mixed Signals is the first V2X dataset that provides heterogeneous CAV LiDAR configurations and features a left-handing driving country, Australia. The dataset includes 45.1k point clouds and 240.6k bounding boxes, collected from three CAVs equipped with two types of LiDAR sensors, along with a roadside unit with two LiDARs. It captures a diverse range of traffic participants across 10 different classes, including 4 vulnerable road user categories. Furthermore, to our best knowledge, Mixed Signals is the highest quality large-scale V2X dataset publicly available, providing precisely aligned and consistently annotated data across both time and different viewpoints. We emphasize that our dataset is ready-to-use; the data and annotations are available on the dataset website (https://mixedsignalsdataset.cs.cornell.edu/), along with the corresponding video visualization showcasing their quality. To summarize, our contributions are as follows:
  • To the best of our knowledge, we are the first real-world V2X dataset that encompasses heterogeneous CAV LiDAR setups and left-hand traffic scenarios.
2 Related Works
@@
V2V4Real [42] USA Right 0 2 20.0 240.0 5 0
V2X-Real [37] USA Right 2 2 33.0 1200.0 10 2
@@
Mixed Signals AUS Left 2 3 45.1 240.6 10 4
Table 1: Comparison of Mixed Signals and existing V2X datasets. To our best knowledge, Mixed Signals is the first dataset to include heterogeneous CAV LiDAR configurations, and also the first one that is collected in a left-hand driving country. It captures complex, real-world traffic scenarios and features a diverse range of traffic participants. Those marked with $\ddagger$ are not publicly accessible anywhere on earth.
@@
Vehicle-to-Everything Communication. One of V2X’s objectives is to enhance the perception capabilities of CAVs, facilitating their deployment in urban environments. These areas usually have a high presence of Vulnerable Road Users (VRUs) which are people not inside vehicles [32]. Despite this, VRUs are under represented in prior works. The three synthetic datasets made with CARLA [10] and the real-world dataset V2V4Real [42] do not have VRUs. DAIR-V2X-C [44] and its extension V2X-Seq (SPD) [45] provide annotations for 4 VRU classes (pedestrian, bicyclist, tricyclist, and motorcyclist). However, the absence of details on class distribution in their publications make it hard to judge their coverage VRUs. In addition, the fact that the access to these two datasets from outside of China is conditional reduces their applicability. TUMTrafV2X [47] annotates 3 VRU classes including pedestrian, bicycle, and motorcycle, which together account for 24.6% of the total annotations. Such under representation leads to the overlook of VRUs detection in several works on collaborative perception methods [36, 22, 41, 40].
Real World Vehicle-to-Everything Datasets. The recent V2X-Real [37] has a large number pedestrian annotations, which is higher than annotations of the class car, and 3 other VRU classes (scooter, motorcycle, and bicycle). A drawback of this dataset to the VRU detection is that their benchmark only accounts for pedestrian. Our dataset contains the highest number of VRU classes, including pedestrian, bicycle, portable personal mobility, and motorcycle. More importantly, these classes account for 50.3% of our dataset’s total bounding boxes. Instead of selecting certain VRU classes for benchmarking, we group 4 VRU classes into 2 detection classes as in Section 3.4 to provide a better understanding of how different collaboration methods perform in detecting VRUs. We provide a detailed comparison of our dataset, Mixed Signals, with prior works in Table 1.
@@ 3 Mixed Signals Dataset
In this section, we describe the data collection process of the Mixed Signals dataset. We have included a devkit and a sample subset of our dataset in the supplementary materials, to allow interested readers to explore it in more detail.
@@ 3.1 Hardware
Vehicles. The three vehicles included two small electric vehicles (EVs) and one urban vehicle, each equipped with OS1 128-beam LiDARs, as shown in Figure 2. The LiDAR on the urban vehicle is located horizontally with respect to the ground, while for the EV, the LiDAR is tilted downwards 15 degrees. We transformed the EVs’ point cloud to have a horizontal reference frame as shown in 2(a). Although all the vehicles are equipped with the same type of LiDAR sensor, their configurations differ in terms of sensor position and orientation. This variation introduces additional complexity, creating a domain gap between the data collected from different vehicles.
(a) Electric vehicle (EV) with a OS1-128 beams LiDAR.
(b) Urban vehicle with a OS1-128 beams LiDAR.
(c) Roadside Unit (RSU) hardware.
@@ 3.2 Data Acquisition
3.2.1 Location
The data collection took place at the intersection between Abercrombie Street and Myrtle Street (Sydney, Austria), where the roadside unit is located. The vehicles recorded LiDAR data for two hours during peak rush hour. Throughout this period, the three vehicles repeatedly passed through the intersection. This allowed them to capture interactions between the vehicles and other agents on the road, such as pedestrians, cyclists, and other vehicles.
@@ 3.2.2 Synchronization and Localization
Synchronization and localization are crucial for cross-sensor point cloud alignment. Our dataset employs proven robotic techniques to achieve precise sensor synchronization and agent localization, resulting in superior point cloud alignment compared to previous V2X datasets (Figure 4). We describe the details below.
Synchronization refers to the temporal alignment of data streams, ensuring that synchronized sensors capture the same events simultaneously within their overlapping fields of view (FOV). This is especially important in dynamic environments, as any introduced time shifts can lead to positional inconsistencies, resulting in multiple detections of the same object.
We use GPS time to timestamp point clouds captured by our LiDARs at a frequency of 10 Hz. Rotational LiDARs continuously scan the environment in 360 degrees, thus, different portions of the surroundings are captured at slightly different times during a full rotation. When vehicles are in motion, their positions and orientations change dynamically between LiDARs sweeps. This means that even if two vehicles are GPS-synchronized, the data they collect will represent slightly different moments in time and from different spatial perspectives. We defined data samples by setting a time window to match the closest available timestamps from each LiDARs. A maximum timestamp mismatch of 50 milliseconds between point clouds was set to achieve minimal spatial discrepancies.
Localization is one of the most critical tasks for CAV, estimating their position relative to a global reference frame. One of the most commonly used sensors for localization is the Global Navigation Satellite System (GNSS). GNSS offers access to a satellite constellation that provides global positioning via triangulation. However, despite its widespread use, GNSS has several drawbacks, particularly in urban environments. Its accuracy can be reduced in urban canyons, where tall buildings block or reflect signals, leading to degraded positioning accuracy.
To overcome this problem, we use dense and accurate point cloud maps [31] as references for our localization algorithm. Both the vehicles and the roadside units are localized within a common reference frame, referred to as the $map\_frame$, which serves as the origin of our map. The localization algorithm employs a scan-matching technique [3] to estimate the vehicles’ poses within this map, achieving a maximum positioning error of 15 cm and a heading error of 0.4 degrees. This allows for consistent spatial alignment between the vehicles and the roadside infrastructure. The vehicles’ localization estimates their positions within the $map\_frame$, while the roadside unit is static.
@@ 3.2.3 Scene Selection
In total, 37 scenes were carefully selected for inclusion in the dataset due to their rich diversity of vehicles, pedestrians, and cyclists. The primary goal was to capture various vehicles and vulnerable road users. These scenes encompass a broad spectrum of interactions, including between different types of vehicles and between vehicles and vulnerable road users. The selected scenes feature intersections of the FOV of the LiDARs of 3 vehicles and the RSU.
Among 37 scenes of our dataset, we select 33 scenes for training and 4 scenes of testing. The size of the training set and test set are, respectively, 9553 and 1164 data samples. Our selection ensures that there is not any temporal overlapping between the training set and test set and among scenes of the test set.
@@ 3.3 Dataset Annotation
The task of 3D object detection for autonomous vehicles requires annotations in the form of 3D bounding boxes, usually parameterized by the center location, three dimensions (length, width, height), and rotation (represented as a quaternion). To generate such annotations for each data sample, we first aggregate the point clouds of every agent in the coordinate of the roadside unit’s top (TOP) LiDAR. Then, professional annotators from FlipSideAI [34] employ the SegmentsAI [35] annotation tool to label objects and localize them with 3D bounding box. Classes labeled belong to 10 categories, consisting of: car, truck, pedestrians, bus, electric vehicle, trailer, motorcycle/bike, bicycle, portable personal mobility, and emergency vehicle. Figure 1 depicts the annotations applied to the dataset, where each object is enclosed within a cuboid. Our annotation process involved cycles of monitoring, reviewing, and adjusting labels to meet defined quality objectives. This allows Mixed Signals dataset to extend the quality of the pioneering datasets in the field, which are generally labeled by lay annotators, as shown in Figure 5. Here, we reproject the bounding box of a vehicle, as observed from other sensors, back onto its coordinate frame to visualize label consistency. Details of the class descriptions and labeling instructions are presented in supplementary materials.
Figure 5: Label quality comparison of Mixed Signals and existing datasets. Labels in Mixed Signals are highly consistent across different time steps and viewpoints.
The annotation process for this multi-sensor dataset involves handling joint scenes and synchronization discrepancies between sensors. Due to time synchronization, fast-moving objects might appear slightly offset across the data collected from different sensors. To address these discrepancies, annotators were instructed to prioritize the roadside unit point cloud for bounding box creation, following a set hierarchy. When there is a mismatch, bounding boxes should be aligned with the point cloud in the following order: roadside unit, EVs, and the urban vehicle. For example, if there is a difference between the roadside unit and the vehicles’ point cloud, the bounding box should only be fitted around the roadside unit points. This ensures consistency in object localization across frames despite synchronization lags. While agents in our dataset are synchronized at 10 Hz, we sample keyframes at 1 Hz for manual annotation. To obtain annotations in a non-key frame, we linearly interpolate the pose of annotations of its closest preceding and succeeding keyframes based on their timestamp.
@@
The Mixed Signals dataset categories road agents in different vehicle types and pedestrians. Categories such as “Car” and “Truck” encompass common passenger and large transport vehicles, while “Emergency Vehicle” covers ambulances, fire trucks, and police cars, highlighting their importance in urban scenarios. “Bus” labels are designated for large passenger vehicles typically used in public transportation. The dataset also distinguishes between “Motorcycle” and “Motorized Bike,” and “Portable Personal Mobility Vehicle,” which includes modern personal transport devices like electric scooters and hoverboards. Traditional “Bicycle” labels account for both standard and electric bikes. Labels for “Electric Vehicle” and “Trailer” ensure that smaller, often data-collection vehicles and towable units are accurately represented. Finally, we labeled humans as “Pedestrians”. Table S1 provides the detailed definition of each category.
Figure 6: Distribution of LiDAR intensities from RSU TOP, DOME, and Laser car sensors. Each sensor shows different intensity ranges and distributions. EV-1 and EV-2 LiDAR sensors do not have intensity readings.

### v2

@@ Abstract
Vehicle-to-everything (V2X) collaborative perception has emerged as a promising solution to address the limitations of single-vehicle perception systems. However, existing V2X datasets are limited in scope, diversity, and quality. To address these gaps, we present Mixed Signals, a comprehensive V2X dataset featuring 45.1k point clouds and 240.6k bounding boxes collected from three connected autonomous vehicles (CAVs) equipped with two different configurations of LiDAR sensors, plus a roadside unit with dual LiDARs. Our dataset provides point clouds and bounding box annotations across 10 classes, ensuring reliable data for perception training. We provide detailed statistical analysis on the quality of our dataset and extensively benchmark existing V2X methods on it. Mixed Signals is ready-to-use, with precise alignment and consistent annotations across time and viewpoints. We hope our work advances research in the emerging, impactful field of V2X perception. Dataset details at https://mixedsignalsdataset.cs.cornell.edu/.
Figure 1: Six samples of the scenes in the dataset. The different agent’s LiDAR point clouds are colored as follows: electric vehicle-001 (EV-1) in purple, electric vehicle-002 (EV-2) in red, urban vehicle (Laser) in yellow, and the RSU DOME and TOP LiDARs in green and blue, respectively. We also draw the annotated bounding boxes within the scene. Best viewed in color.
@@ 1 Introduction
In recent years, driver assistance [19, 29] and autonomous driving [47, 1] technologies have advanced significantly, equipping vehicles with promising capabilities in perception [20, 44], planning [13, 15], and control [2, 8]. Most of these developments focus on single autonomous vehicle scenarios. Despite the advancements, such settings still face challenges in complex or unpredictable situations [42]. For instance, important traffic participants can be occluded from view, or sensors can fail unexpectedly. As autonomous vehicle deployment increases, new possibilities emerge to address these issues: multiple vehicles can communicate with each other and nearby infrastructure, enabling each vehicle to reliably detect road users even when its own sensors miss them by leveraging shared information. This approach is commonly referred as vehicle-to-everything (V2X) collaborative perception.
@@
V2V4Real [43] Height USA Right 0 2 20.0 240.0 5 0
V2X-Real [38] Height USA Right 2 2 33.0 1200.0 10 2
@@
Mixed Signals Height, Tilt AUS Left 2 3 45.1 240.6 10 4
Table 1: Comparison of Mixed Signals and existing V2X datasets. To our best knowledge, Mixed Signals is the first dataset to include heterogeneous CAV LiDAR configurations, and also the first one that is collected in a left-hand driving country. It captures complex, real-world traffic scenarios and features a diverse range of traffic participants. Those marked with $\ddagger$ are valuable datasets, but are only accessible from certain geographical regions.
While single-vehicle perception datasets are abundant across diverse driving conditions [11, 18, 33, 4, 5, 28, 9, 39, 16, 27, 30], real-world V2X datasets remain limited in availability, diversity, and quality. Only a handful of publicly available V2X datasets exist [43, 38, 48, 24], with some of them accessible only within specific geographical regions [45, 46, 12, 26]. These data are collected exclusively from three right-hand traffic locations, overlooking the unique traffic dynamics in left-hand traffic countries which make up about a third of the world [40]. Furthermore, as collaborative perception becomes more widespread, it is valuable for vehicles equipped with different sensor configurations to communicate. However, in prior datasets, the connected autonomous vehicles (CAVs) share identical or very similar LiDAR configurations. Finally, as the V2X setting involves multiple agents and sensors, data collection and alignment present additional challenges. Often times, difficulty with pose estimation and faulty localization systems result in poor alignment (Figure 4). Such inaccuracies can lead to suboptimal performance for detector training [41].
To address these limitations, we introduce the Mixed Signals dataset, designed to support diverse real-world V2X research scenarios with clean, high-quality data. Notably, Mixed Signals is the first V2X dataset that provides heterogeneous CAV LiDAR configurations in both position and orientation, and features a left-handing traffic country, Australia. The dataset includes 45.1k point clouds and 240.6k bounding boxes, collected from three CAVs equipped with two configurations of LiDAR sensors, along with a roadside unit with two LiDARs. It captures a diverse range of traffic participants across 10 different classes, including 4 vulnerable road user categories. Furthermore, compared to existing datasets, Mixed Signals offers significantly more precise alignment and consistent annotations across time and viewpoints. We emphasize that our dataset is ready-to-use; a subset is provided in the supplementary materials, along with the corresponding video visualization showcasing the quality of our collected data and annotations. To summarize, our contributions are:
  • To the best of our knowledge, we are the first real-world V2X dataset that encompasses CAV LiDAR configurations that differ in both position and orientation, as well as left-hand traffic scenarios.
2 Related Works
@@
(a) Electric vehicle (EV).
(b) Urban vehicle with a OS1-128 beams LiDAR.
(c) Roadside Unit (RSU).
Vehicle-to-Everything Communication. One of V2X’s objectives is to enhance the perception capabilities of CAVs, facilitating their deployment in urban environments. These areas usually have a high presence of Vulnerable Road Users (VRUs), which are defined as people not inside vehicles [32]. Despite this, VRUs are under represented in prior works. The three synthetic datasets made with CARLA [10] and the real-world dataset V2V4Real [43] do not have VRUs. DAIR-V2X-C [45] and its extension V2X-Seq (SPD) [46] provide annotations for 4 VRU classes (pedestrian, bicyclist, tricyclist, and motorcyclist). However, the absence of details on class distribution in their publications make it hard to judge their VRU coverage. Additionally, restricted access to these datasets outside China limits their usability. TUMTrafV2X [48] annotates 3 VRU classes including pedestrian, bicycle, and motorcycle, which together account for only 24.6% of the total annotations. Such underrepresentation causes VRU detection to be overlooked in several collaborative perception studies [36, 22, 42, 41].
Real World Vehicle-to-Everything Datasets. The recent V2X-Real [38] has a large number pedestrian annotations, which is higher than annotations of the class car, and 3 other VRU classes (scooter, motorcycle, and bicycle). A drawback of this dataset for VRU detection evaluation is that its benchmark only accounts for pedestrians. Our dataset contains the highest number of VRU classes, including pedestrian, bicycle, portable personal mobility, and motorcycle. More importantly, these classes account for 50.3% of our dataset’s total bounding boxes. Instead of selecting certain VRU classes for benchmarking, we group 4 VRU classes into 2 detection classes as in subsection 3.4 to provide a better understanding of how different collaboration methods perform in detecting VRUs. We provide a detailed comparison of our dataset, Mixed Signals, with prior works in Table 1.
@@ 3 Mixed Signals Dataset
In this section, we describe the data collection process of the Mixed Signals dataset. We provide a devkit and our full dataset for download on our website: https://mixedsignalsdataset.cs.cornell.edu/.
@@ 3.1 Hardware
Vehicles. The three vehicles included two small electric vehicles (EVs) and one urban vehicle, each equipped with OS1 128-beam LiDARs, as shown in Figure 2. The LiDAR on the urban vehicle is located horizontally with respect to the ground, while for the EV, the LiDAR is tilted downwards 15 degrees. We transformed both EVs’ point clouds to have a horizontal reference frame as shown in 2(a). Although all the vehicles are equipped with the same type of LiDAR sensor, their configurations differ in terms of sensor position and orientation. This variation introduces additional complexity, creating a domain gap between the data collected from different vehicles.
@@ 3.2 Data Acquisition
The data collection took place at the intersection between Abercrombie Street and Myrtle Street in Sydney, Austria, where the roadside unit is located. The vehicles recorded LiDAR data for two hours during peak rush hour. Throughout this period, the three vehicles repeatedly passed through the intersection. This allowed them to capture interactions between the vehicles and other agents on the road, such as pedestrians, cyclists, and other vehicles.
@@ 3.2.1 Synchronization and Localization
Synchronization and localization are crucial for cross-sensor point cloud alignment. Our dataset employs proven techniques from robotics to achieve precise sensor synchronization and agent localization. The end result is superior point cloud alignment compared to previous V2X datasets (Figure 4). We describe the details below.
Synchronization refers to the temporal alignment of data streams, ensuring that synchronized sensors capture the same events simultaneously within their overlapping fields of view (FOV). We use GPS time to timestamp point clouds captured by our LiDARs at a frequency of 10 Hz. Even if two vehicles are GPS-synchronized, cross-sensor synchronization still needs to be considered. For example, since the LiDAR scans the environment in a rotating fashion, the data collected at different spatial locations are captured at slightly different moments. We defined data samples by setting a time window to match the closest available timestamps from each LiDARs. A maximum timestamp mismatch of 50 milliseconds between point clouds was set to achieve minimal spatial discrepancies. For additional details, refer to Appendix C.1.
Localization, i.e., estimating vehicle position relative to a global reference frame, is one of the most critical tasks for CAV. To overcome inherent problems of Global Navigation Satellite System (GNSS) in urban environments, we use dense and accurate point cloud maps [31] as references for our localization algorithm. Both the vehicles and the roadside units are localized within a common reference frame, referred to as the $map\_frame$, which serves as the origin of our map. The localization algorithm employs a scan-matching technique [3] to estimate the vehicles’ poses within this map, achieving a maximum positioning error of 15 cm and a heading error of 0.4 degrees. This allows for consistent spatial alignment between the vehicles and the roadside infrastructure. The vehicles’ localization estimates their positions within the $map\_frame$, while the roadside unit is static. We leave details about map construction and usage in Appendix C.2.
@@ 3.2.2 Scene Selection
In total, 37 scenes –each consisting of a 30-second snippet– were carefully selected for inclusion in the dataset due to their rich diversity of vehicles, pedestrians, and cyclists. The primary goal was to capture various vehicles and vulnerable road users. These scenes encompass a broad spectrum of interactions, including between different types of vehicles and between vehicles and vulnerable road users. The selected scenes feature intersections of the FOV of the LiDARs of 3 vehicles and the RSU. Among 37 scenes of our dataset, we select 33 scenes for training and 4 scenes of testing. The size of the training set and test set are 9553 and 1164 data samples, respectively. Our selection ensures that there is no temporal overlap between the training set and test set and among scenes of the test set.
@@ 3.3 Dataset Annotation
The task of 3D object detection for autonomous vehicles requires annotations in the form of 3D bounding boxes, usually parameterized by the center location, three dimensions (length, width, height), and rotation (represented as a quaternion). To generate such annotations for each data sample, we first aggregate the point clouds of every agent in the coordinate of the roadside unit’s top (TOP) LiDAR to focus the annotators’ attention to the intersection of interest. Then, professional annotators from FlipSideAI [34] employ the SegmentsAI [35] annotation tool to label objects and localize them with a 3D bounding box. Classes labeled belong to 10 categories, consisting of: car, truck, pedestrians, bus, electric vehicle, trailer, motorcycle/bike, bicycle, portable personal mobility, and emergency vehicle. Figure 1 depicts the annotations applied to the dataset, where each object is enclosed within a cuboid.
Annotations. Our annotation process involved cycles of monitoring, reviewing, and adjusting labels to meet defined quality objectives. This allows Mixed Signals dataset to extend the quality of the pioneering datasets in the field, which are generally labeled by lay annotators, as shown in Figure 5. Here, we reproject the bounding box of a vehicle, as observed from other sensors, back onto its coordinate frame to visualize label consistency. Details of the class descriptions and labeling instructions are presented in Appendix Sec.  B. While agents in our dataset are synchronized at 10 Hz, we sample keyframes at 1 Hz for manual annotation. To obtain annotations in a non-key frame, we linearly interpolate the pose of annotations of its closest preceding and succeeding keyframes based on their timestamp.
Figure 5: Label quality of Mixed Signals and existing datasets. We aggregate labels of an object across a entire snippet. Labels in Mixed Signals are consistent across time steps and viewpoints.
@@
The Mixed Signals dataset categories consist of road agents in 10 categories of vehicle types and pedestrians including: Car, Truck, Emergency Vehicle, Bus, Motorcycle, Motorized Bike, Portable Personal Mobility Vehicle, (traditional) Bicycle, Electric Vehicle, Trailer, and Pedestrian. Detailed definition of each category can be found in the appendix.